OpenStudy (anonymous):

what is the solution to the system of equations 4x+y=2 x-y=3

OpenStudy (sleepyjess):

Welcome to OpenStudy! We can solve this by elimination or substitution, it's up to you :)

OpenStudy (anonymous):

yeah

OpenStudy (sleepyjess):

Which one do you want to use?

OpenStudy (anonymous):

4x+y=2

OpenStudy (sleepyjess):

We need to solve both equations by substitution or elimination

OpenStudy (anonymous):

substitution

OpenStudy (sleepyjess):

Okay :) So we need to get x or by by itself in either equation, I think x-y = 3 will be easiest to work with, what do you think?

OpenStudy (anonymous):

yes

OpenStudy (sleepyjess):

Okay, so let's get x by itself since it's already positive :) do you know how to do that?

OpenStudy (anonymous):

take it out ?

OpenStudy (sleepyjess):

We need to move -y to the right side of the equation

OpenStudy (anonymous):

Dont We Need To Put Those Together

OpenStudy (sleepyjess):

Put what together?

OpenStudy (anonymous):

4x+y=2 x-y=3

OpenStudy (sleepyjess):

That is elimination

OpenStudy (anonymous):

Which Would Be Easier

OpenStudy (sleepyjess):

It really doesn't matter with this one, both ways will be about the same, but probably elimination if you know how to do it